# Improving Molecule Generation and Drug Discovery with a Knowledge-enhanced Generative Model

### v1

@@ 5.1 Property Inference Network $P_{\phi}(\textbf{G})$
5.2 Conditional Diffusion Training
The stochastic forward process described in Eq. (1) can be used for generation by solving it’s reverse-time version:
@@ 5.2 Conditional Diffusion Training
where $\alpha_{\theta}$ is a scaling factor and $Z_{\theta}$ is the partition function. This procedure has a tractable complexity, supported by the analysis in Section 4.
6 Experiments
We perform individual evaluations of each component, followed by an end-to-end study of K-DReAM in synthesizing novel drug candidates. We compare unconditional generation against variants of generative models that use GANs (Martinkus et al., 2022), diffusion (Jo et al., 2022a; Vignac et al., 2023), autoregressive methods (Kong et al., 2023), and normalizing flows (Shi et al., 2019; Zang & Wang, 2020).
6.1 Evaluating Unconditional Molecular Generation
Experimental Setup. The Quantum Machines 9 (QM9) (Ramakrishnan et al., 2014) and ZINC (Irwin et al., 2012) datasets contain 134k and 250k chemical structures and are designed to aid in the exploration of chemical space. We follow previous work (Kong et al., 2023) and perform unconditional generation of 10,000 molecular structures. We report the percentage of valid and unique molecules, along with the novelty as defined by (Jin et al., 2020b) and the Frechet ChemNet Distance (FCD) (Preuer et al., 2018), which evaluates the similarity between the generated and training datasets using the activations of the penultimate layer of the ChemNet. The reward function (Eq. (3)) is used to fine-tune K-DReAM for maximizing the validity and novelty of the molecules.
Results. Tables 1 and 2 demonstrate that K-DReAM achieves state-of-the-art results on all metrics except the FCD on the QM9 dataset. We omit reporting uniqueness on ZINC since all models were $\gtrapprox 99.99\%$ unique. Previous results sourced from (Kong et al., 2023). K-DReAM generates valid molecules at rates of 99.25% and 98.29% on QM9 and ZINC, respectively as a result of the fine-tuning procedure.
@@ 6.1 Evaluating Unconditional Molecular Generation
K-DReAM (w/o RL) 94.20 2.21 97.64 95.63
K-DReAM 99.25 1.54 98.62 99.74
Table 1: Generation results on the QM9 dataset. (Best, Second)
@@
K-DReAM (w/o RL) 96.17 17.77 97.64
K-DReAM 98.29 14.56 99.95
Table 2: Generation results on ZINC. (Best, Second)
6.2 Drug-based Knowledge Graph Embeddings
Experimental Setup. (Bonner et al., 2022) devise a procedure to assess how well biomedical KGE models can predict missing links in knowledge graphs, standard evaluation employs knowledge graph metrics like Mean Reciprocal Rank and Hits@k. These metrics measure the model’s ability to rank true triples higher than corrupted ones in link prediction tasks. Models are evaluated on predicting both head and tail entities by corrupting each in turn. Robustness is examined by testing across multiple random parameter initializations and dataset splits. Performance is evaluated on real-world KGs - BioKG (Walsh et al., 2020) and Hetionet (Himmelstein et al., 2017). The following metrics are used for comparison: (i) Mean Reciprocal Rank (MRR) (ii) Hits@1 and Hits@10 (iii) Adjusted Mean Rank (Berrendorf et al., 2020). Hits@k is computed using the following procedure: For each corrupted triple (with head or tail entity removed) in the test set, the model ranks all possible candidate entities to complete the triple. A “hit” is when the true missing entity is ranked within the top $k$ predictions. Hits@k calculates the proportion of test triples where the true entity was ranked in the top $k$.
Results. Table 3 presents the mean performance over 10 random seeds of all models and both datasets as measured by the above metrics. Only the random seed used to initialize the model parameters is changed. The mean score of K-DReAM outperforms all metrics except the AMR score on BioKG, showing the improved quality of embeddings due to the introduction of domain constraints. We compare against ComplEx (Trouillon et al., 2017), DistMult (Zhang et al., 2018), RotatE (Sun et al., 2019), and TransE (Lin et al., 2015).
@@ 6.2 Drug-based Knowledge Graph Embeddings
K-DReAM 0.025$\pm$0.010 0.136$\pm$0.010 0.065$\pm$0.000 0.28$\pm$0.001
Table 3: Evaluation of KGE Model. Previous results sourced from (Bonner et al., 2022).
@@
REINVENT (Olivecrona et al., 2017) 0-8.702 ($\pm$ 0.523) -7.205 ($\pm$ 0.264) 0-8.770 ($\pm$ 0.316) 0-8.392 ($\pm$ 0.400) 0-8.165 ($\pm$ 0.277)
MORLD (Jeon & Kim, 2020) 0-7.532 ($\pm$ 0.260) -6.263 ($\pm$ 0.165) 0-7.869 ($\pm$ 0.650) 0-8.040 ($\pm$ 0.337) 0-7.816 ($\pm$ 0.133)
HierVAE (Jin et al., 2020a) 0-9.487 ($\pm$ 0.278) -6.812 ($\pm$ 0.274) 0-8.081 ($\pm$ 0.252) 0-8.978 ($\pm$ 0.525) 0-8.285 ($\pm$ 0.370)
FREED (Yang et al., 2021) -10.427 ($\pm$ 0.177) -8.297 ($\pm$ 0.094) -10.425 ($\pm$ 0.331) -10.325 ($\pm$ 0.164) 0-9.624 ($\pm$ 0.102)
@@
K-DReAM (w/o KGE conditioning) -9.416 ($\pm$ 0.463) -7.920 ($\pm$ 0.321) -9.679 ($\pm$ 0.345) -9.465 ($\pm$ 0.098) -9.782 ($\pm$ 0.303)
K-DReAM -11.475 ($\pm$ 0.410) -8.270 ($\pm$ 0.211) -11.419 ($\pm$ 0.072) -11.364 ($\pm$ 0.035) -10.636 ($\pm$ 0.211)
@@
Table 4: Top 5% docking score (kcal/mol). The results are the means and the standard deviations of 5 runs. Previous results sourced from (Lee et al., 2022).
6.3 Knowledge-enhanced Drug Discovery: Novelty and Protein Targeting
@@
Experimental Setup. K-DReAM generates 3,000 drug candidates with the reward function (Eq. (3)) calibrated to optimize synthesizability, drug-likeness and binding affinity to a target protein. Following (Lee et al., 2022), we choose the targets: parp1, fa7, 5ht1b, braf, jak2. We set the chemical property $C$ to emphasize novelty, introducing a penalty for a similarity score $>0.4$ with the training dataset. The knowledge embeddings are generated by asking the model to complete the triple (_____, targets, protein) for all five targets, searching over the molecule space for the highest scoring embeddings. These embeddings are used for conditional generation as desctribed in Section 5.
@@
Results. Table 4 reports the average Docking Score (DS) of the top 5% of generated molecules. K-DReAM is able to generate novel molecules with the best docking scores among the models we compare against, with previous results sourced from (Lee et al., 2022). The procedure for calculating these scores is described in Appendix A.
6.4 Ablation Study: Generation without Reinforcement Learning
@@
We perform an ablation study to examine the effect of fine-tuning K-DReAM with the reward function (Eq. (3)) by comparing the performance before the fine-tuning procedure in Tables 1 and 2. We find that even without DDPO, K-DReAM has a similar performance as other diffusion- and SDE-based systems. DDPO improves the validity and novelty to $\geq 99\%$, an improvement of 4-5% points over the non fine-tuned performance.
6.5 Ablation Study: Generation without Knowledge Enhancement
@@
We validate the impact of integrating KGEs into the model by running the target protein experiments with a version of K-DReAM that is fine-tuned using DDPO but not trained for conditional generation. Table 4 shows that K-DReAM without KGE information obtains molecules that do not score as well in the docking metric. In most cases, the docking score is similar to that of GDSS, which is in line with our expectations, since our base model is derived from a similar framework. The DDPO fine-tuning is identical in both cases. We thus demonstrate that the incorporation of knowledge through embeddings leads to a measurable improvement in targeted drug parameters. We can also consider a comparison between K-DReAM and RotatE in Table 3 as an ablation study, since our model is an improved version of RotatE with the modifications in Section 4.
7 Conclusion
@@
We enhance molecular generation and drug discovery models by incorporating information from biomedical knowledge graphs into the generative process. This framework, called K-DReAM, demonstrates: 1) Performance on unconditional molecular generation on par with state-of-the-art generative models, 2) Improvements in Knowledge Graph Embedding (KGE) metrics by introducing domain constraints and reformulating the scoring process, and 3) Enhancement with KGEs improves targeted generation, evaluated with docking scores for five proteins.
@@
8 Acknowledgements
This work uses the Anvil supercomputer (Song et al., 2022) at Purdue University through allocation CIS230228 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The authors gratefully acknowledge the Walther Cancer Foundation and support from the Purdue University Institute for Cancer Research, P30CA023168
Appendix A Implementation Details
@@
A.1 Knowledge graphs
PyKEEN (Ali et al., 2021) (built using Pytorch) is used to implement the Knowledge Graph Embedding Model (KGEM). The knowledge graph dataloaders for BioKG and Hetionet in PyKEEN are used with a custom defined model based on the formulation of (Loconte et al., 2023).
A.2 Docking scores

### v2

@@ V-A Property Inference Network $P_{\phi}(\textbf{G})$
V-B Conditional Diffusion Training
The stochastic forward process described in Eq. (1) can be used for generation by solving it’s reverse-time version (Fig 2):
@@ V-B Conditional Diffusion Training
where $\alpha_{\theta}$ is a scaling factor and $Z_{\theta}$ is the partition function. This procedure has a tractable complexity, supported by the analysis in Section IV.
VI Experiments
We perform individual evaluations of each component, followed by an end-to-end study of KARL in synthesizing novel drug candidates. We compare unconditional generation against variants of generative models that use GANs [17], diffusion [20, 21], autoregressive methods [22], and normalizing flows [18, 19].
VI-A Evaluating Unconditional Molecular Generation
Experimental Setup. The Quantum Machines 9 (QM9) [45] and ZINC [46] datasets contain 134k and 250k chemical structures and are designed to aid in the exploration of chemical space. We follow previous work [22] and perform unconditional generation of 10,000 molecular structures. We report the percentage of valid and unique molecules, along with the novelty as defined by [47] and the Frechet ChemNet Distance (FCD) [48], which evaluates the similarity between the generated and training datasets using the activations of the penultimate layer of the ChemNet. The reward function (Eq. (3)) is used to fine-tune KARL for maximizing the validity and novelty of the molecules.
Results. Tables I and II demonstrate that KARL achieves state-of-the-art results on all metrics except the FCD on the QM9 dataset. We omit reporting uniqueness on ZINC since all models were $\gtrapprox 99.99\%$ unique. Previous results sourced from [22]. KARL generates valid molecules at rates of 99.25% and 98.29% on QM9 and ZINC, respectively as a result of the fine-tuning procedure.
@@ VI-A Evaluating Unconditional Molecular Generation
KARL (w/o RL) 94.20 2.21 97.64 95.63
KARL 99.25 1.54 98.62 99.74
TABLE I: Generation results on the QM9 dataset. (Best, Second)
@@
KARL (w/o RL) 96.17 17.77 97.64
KARL 98.29 14.56 99.95
TABLE II: Generation results on ZINC. (Best, Second)
VI-B Drug-based Knowledge Graph Embeddings
Experimental Setup. [49] devise a procedure to assess how well biomedical KGE models can predict missing links in knowledge graphs, standard evaluation employs knowledge graph metrics like Mean Reciprocal Rank and Hits@k. These metrics measure the model’s ability to rank true triples higher than corrupted ones in link prediction tasks. Models are evaluated on predicting both head and tail entities by corrupting each in turn. Robustness is examined by testing across multiple random parameter initializations and dataset splits. Performance is evaluated on real-world KGs - BioKG [37] and Hetionet [36]. The following metrics are used for comparison: (i) Mean Reciprocal Rank (MRR) (ii) Hits@1 and Hits@10 (iii) Adjusted Mean Rank [50]. Hits@k is computed using the following procedure: For each corrupted triple (with head or tail entity removed) in the test set, the model ranks all possible candidate entities to complete the triple. A “hit” is when the true missing entity is ranked within the top $k$ predictions. Hits@k calculates the proportion of test triples where the true entity was ranked in the top $k$.
Results. Table III presents the mean performance over 10 random seeds of all models and both datasets as measured by the above metrics. Only the random seed used to initialize the model parameters is changed. The mean score of KARL outperforms all metrics except the AMR score on BioKG, showing the improved quality of embeddings due to the introduction of domain constraints. We compare against ComplEx [51], DistMult [52], RotatE [43], and TransE [38].
@@ VI-B Drug-based Knowledge Graph Embeddings
KARL 0.025$\pm$0.010 0.136$\pm$0.010 0.065$\pm$0.000 0.28$\pm$0.001
TABLE III: Evaluation of KGE Model. Previous results sourced from [49].
@@
REINVENT [53] 0-8.702 ($\pm$ 0.523) -7.205 ($\pm$ 0.264) 0-8.770 ($\pm$ 0.316) 0-8.392 ($\pm$ 0.400) 0-8.165 ($\pm$ 0.277)
MORLD [54] 0-7.532 ($\pm$ 0.260) -6.263 ($\pm$ 0.165) 0-7.869 ($\pm$ 0.650) 0-8.040 ($\pm$ 0.337) 0-7.816 ($\pm$ 0.133)
HierVAE [55] 0-9.487 ($\pm$ 0.278) -6.812 ($\pm$ 0.274) 0-8.081 ($\pm$ 0.252) 0-8.978 ($\pm$ 0.525) 0-8.285 ($\pm$ 0.370)
FREED [56] -10.427 ($\pm$ 0.177) -8.297 ($\pm$ 0.094) -10.425 ($\pm$ 0.331) -10.325 ($\pm$ 0.164) 0-9.624 ($\pm$ 0.102)
@@
KARL (w/o KGE conditioning) -9.416 ($\pm$ 0.463) -7.920 ($\pm$ 0.321) -9.679 ($\pm$ 0.345) -9.465 ($\pm$ 0.098) -9.782 ($\pm$ 0.303)
KARL -11.475 ($\pm$ 0.410) -8.270 ($\pm$ 0.211) -11.419 ($\pm$ 0.072) -11.364 ($\pm$ 0.035) -10.636 ($\pm$ 0.211)
@@
TABLE IV: Top 5% docking score (kcal/mol). The results are the means and the standard deviations of 5 runs. Previous results sourced from [33].
VI-C Knowledge-enhanced Drug Discovery: Novelty and Protein Targeting
@@
Experimental Setup. KARL generates 3,000 drug candidates with the reward function (Eq. (3)) calibrated to optimize synthesizability, drug-likeness and binding affinity to a target protein. The knowledge graph used to generate embeddings and train the Property Inference Network is BioKG [37]. Following [33], we choose the targets: parp1, fa7, 5ht1b, braf, jak2. We set the chemical property $C$ to emphasize novelty, introducing a penalty for a similarity score $>0.4$ with the training dataset. The knowledge embeddings are generated by asking the model to complete the triple (_____, targets, protein) for all five targets, searching over the molecule space for the highest scoring embeddings. These embeddings are used for conditional generation as desctribed in Section V.
@@
Results. Table IV reports the average Docking Score (DS) of the top 5% of generated molecules. KARL is able to generate novel molecules with the best docking scores among the models we compare against, with previous results sourced from [33]. The procedure for calculating these scores is described in Appendix B.
VI-D Ablation Study: Generation without Reinforcement Learning
@@
We perform an ablation study to examine the effect of fine-tuning KARL with the reward function (Eq. (3)) by comparing the performance before the fine-tuning procedure in Tables I and II. We find that even without DDPO, KARL has a similar performance as other diffusion- and SDE-based systems. DDPO improves the validity and novelty to $\geq 99\%$, an improvement of 4-5% points over the non fine-tuned performance.
VI-E Ablation Study: Generation without Knowledge Enhancement
@@
We validate the impact of integrating KGEs into the model by running the target protein experiments with a version of KARL that is fine-tuned using DDPO but not trained for conditional generation. Table IV shows that KARL without KGE information obtains molecules that do not score as well in the docking metric. In most cases, the docking score is similar to that of GDSS, which is in line with our expectations, since our base model is derived from a similar framework. The DDPO fine-tuning is identical in both cases. We thus demonstrate that the incorporation of knowledge through embeddings leads to a measurable improvement in targeted drug parameters. We can also consider a comparison between KARL and RotatE in Table III as an ablation study, since our model is an improved version of RotatE with the modifications in Section IV.
VII Conclusion
@@
We enhance molecular generation and drug discovery models by incorporating information from biomedical knowledge graphs into the generative process. This framework, called KARL, demonstrates: 1) Performance on unconditional molecular generation on par with state-of-the-art generative models, 2) Improvements in Knowledge Graph Embedding (KGE) metrics by introducing domain constraints and reformulating the scoring process, and 3) Enhancement with KGEs improves targeted generation, evaluated with docking scores for five proteins.
@@
Appendix A Disclosure of Use of LLM Tools
The authors use the claude.ai model from Anthropic to summarize papers and refine human-written text, and scite.ai to assist with literature review.
Appendix B Implementation Details
@@
B-A Knowledge graphs
PyKEEN [58] (built using Pytorch) is used to implement the Knowledge Graph Embedding Model (KGEM). The knowledge graph dataloaders for BioKG and Hetionet in PyKEEN are used with a custom defined model based on the formulation of [41]. We implement TransE with an embedding dimension of 256. We use the scoring function
$f(h,r,t)=-\|\textbf{e}_{h}+\textbf{e}_{r}-\textbf{e}_{t}\|_{p}$
with $p=1$. We use the Adam optimizer with a learning rate of 0.0001 and a batch size of 16384. The model is trained for 100 epochs. The model is trained using a Stochastic Local Closed World Assumption (SLCWA).
B-B Docking scores
@@
B-C Training the Property Inference Network
The property inference network $P_{\phi}$ learns to estimate Knowledge Graph Embeddings (KGE), given a noised version of the graph, i.e., $P_{\phi}(\textbf{G}_{t})\approx\textbf{c}$. We train this network using a learning schedule starting with the inputs being the original graphs $\textbf{G}_{0}$ and the targets set to their corresponding KGEs c. We then use the forward process $t$ times, where $t<=0.05\times S$, and $S$ is the number of generation steps. These noised graphs are similarly used as inputs with the label still being set to c in order to train the network to approximate to this KGE in the neighborhood of the original graph $\textbf{G}_{0}$.
B-D Adapting Denoising Discrete Policy Optimization (DDPO) for Graphs
We find that the DDPO formulation developed by Black et al. (2023) for images does not adapt well when directly applied to graphs, since the algorithm does not converge as the number of nodes increases. This can mainly be attributed to the variable topology of graphs and the presence of node and edge attributes that lead to an increase in fluctuations.
The MDP for our process is defined as follows:
$s_{t}=(\textbf{G}_{t},T-t),a_{t}=\textbf{G}_{T-t-1}$
$\pi_{\theta}(a_{t}|s_{t}):=p_{\theta}(\textbf{G}_{T-t-1}|\textbf{G}_{T-t})$
$r(s_{t},a_{t})=r(\textbf{G}_{0})\text{ if }t=T;0\text{ otherwise}$
The above framework is generally optimized using the following policy gradient:
$\nabla_{\theta}\mathcal{J}(\theta)=E_{\tau}\left[r(\textbf{G}_{0})\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(\textbf{G}_{t-1}|\textbf{G}_{t})\right]$
We find that this formulation is inadequate and does not converge as the number of nodes increases, and we will include a plot in the paper showing reducing performance as the nodes increase.
Our key improvement is to modify the gradient in the following way:
$\nabla_{\theta}\mathcal{J}(\theta)=E_{\tau}\left[r(\textbf{G}_{0})\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(\textbf{G}_{0}|\textbf{G}_{t})\right]$
Here, $\tau$ is the generation trajectory $\textbf{G}_{0:T}$.
By changing $\nabla_{\theta}\log p_{\theta}(\textbf{G}_{t-1}|\textbf{G}_{t})$ to $\nabla_{\theta}\log p_{\theta}(\textbf{G}_{0}|\textbf{G}_{t})$, we encourage the model to converge as close as possible to $\textbf{G}_{0}$, which has the highest reward at all timesteps. We find that skipping the complexity of considering all the transitions between the intermediate states $\textbf{G}_{t}\to\textbf{G}_{t+1}$ leads to a significantly reduced amount of fluctuations and better convergence.
The parameter is then updated using a learning rate $\eta=10^{–5}$:
$\theta\leftarrow\theta+\eta\cdot\nabla_{\theta}\mathcal{J}(\theta)$
B-E Hyperparamters
For unconditional generation, we use the following hyperparameters:
  • $\kappa_{1}=\kappa_{2}=0$
  • $\kappa_{3}=1$
  • $C=0.7r_{val}+0.2r_{nov}$
For targeted generation, we use the following hyperparameters:
  • $\kappa_{1}=0.4$
  • $\kappa_{2}=0.3$
  • $\kappa_{3}=0.2$
  • $C=0.7r_{val}+0.3r_{nov}$
These are stated in reference to Eq (3) in the main text.
Appendix C Algorithm Tables
C-A Training TransE on BioKG with SLCWA
0:  BioKG dataset $\mathcal{G}=\{(h,r,t)\}$, embedding size $d=256$, learning rate $\alpha=0.005$, margin $\gamma=0.2$, number of epochs $N=100$
0:  Trained TransE embeddings for entities and relations
1:  Initialize:
2:   Entity embeddings $E\in\mathbb{R}^{|E|\times d}$ randomly
3:   Relation embeddings $R\in\mathbb{R}^{|R|\times d}$ randomly
4:  Set Optimizer: Adam with learning rate $\alpha$
5:  for epoch $=1$ to $N$ do
6:     for each triple $(h,r,t)\in\mathcal{G}$ do
7:        // Generate negative samples using SLCWA
8:        Select a corrupted triple $(h^{\prime},r,t)$ or $(h,r,t^{\prime})$ where $h^{\prime}$ or $t^{\prime}$ is sampled locally
9:        // Compute scores
10:        $score_{pos}\leftarrow\|E[h]+R[r]-E[t]\|_{2}$
11:        $score_{neg}\leftarrow\|E[h^{\prime}]+R[r]-E[t^{\prime}]\|_{2}$
12:        // Compute margin-based loss
13:        $loss\leftarrow\max(0,score_{pos}+\gamma-score_{neg})$
14:        // Backpropagate and update embeddings
15:        Compute gradients of $loss$ w.r.t $E[h]$, $E[t]$, $E[h^{\prime}]$, $R[r]$
16:        Update $E[h]$, $E[t]$, $E[h^{\prime}]$, $R[r]$ using Adam optimizer
17:     end for
18:  end for
C-B Score-Based Diffusion Model
0:  Node feature matrix $\mathbf{X}$, Adjacency matrix $\mathbf{A}$, Number of diffusion steps $T$, Score network parameters $\theta$, Guidance scale $\lambda$, Context vector c
0:  Trained score network parameters $\theta$
1:  Initialize score network parameters $\theta$
2:  for epoch = 1 to $N_{\text{epochs}}$ do
3:     for batch in data do
4:        for $t=1$ to $T$ do
5:           Sample noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
6:           Compute noisy features $\mathbf{X}_{t}=\sqrt{\alpha_{t}}\mathbf{X}+\sqrt{1-\alpha_{t}}\epsilon$
7:           Compute score: $\mathbf{s}_{\theta}(\mathbf{X}_{t},\mathbf{A},t)=\nabla_{\textbf{G}_{t}}\log p_{\theta}(\textbf{X}_{t},\textbf{A}_{t})$
8:           Update guidance: $s^{\prime}_{\theta}\leftarrow s_{\theta}+\lambda\nabla_{\textbf{G}_{t}}\log p_{\theta}(\textbf{c}|\textbf{X}_{t},\textbf{A}_{t})$
9:           Compute loss $\mathcal{L}=\|\mathbf{s}^{\prime}_{\theta}(\mathbf{X}_{t},\mathbf{A},t)+\frac{\epsilon}{\sqrt{1-\alpha_{t}}}\|^{2}$
10:        end for
11:        Update $\theta$ using optimizer with gradient $\nabla_{\theta}\mathcal{L}$
12:     end for
13:  end for
14:  Return $\theta$